1. Conditions for price discrimination Aa Aa Price discrimination is the practice of selling the same good at more than one price when the price differences are not justified by cost differences. Evaluate the following statement: Price discrimination is not possible when a good is sold in a competitive market. O False, because competitive firms have market power O False, because competitive firms do not profit maximize by setting marginal revenue equal to marginal cost None of these choices O True, because competitive firms have no market power
Answer #1

1. True, because competitive firms have no market power.

Perfect Competition is a form of market structure in which there is free entry and exit of firms and firms are selling homogeneous and identical products in the market. Firms under this form of market are price takers rather than price makers. Industry determines the equilibrium price from the demand and supply curve intersection. Sellers can sell any unit of commodity at that price and firms does not have any price control over the commodity. If one seller try to charge higher price then it will lose all his customers because all firms are selling similar products in every respect like color, shape, brand, etc.

2. At price of $ 8, quantity demanded in Market A = 6 tickets

Market B = 2 tickets

Total quantity demanded = 6 + 2 = 8 tickets

Total revenue is maximised where MR = 0 i.e. 5 tickets in Market A and 3 tickets in Market B.

Darnell should charge = $ 10 in Market A and $ 6 in Market B.

Total quantity = 5 + 3 = 8 tickets

Tota revenue at same price = $ 8 x 8 tickets = $ 64

Total revenue at different prices = 10 x 5 + 6 x 3 = 50 + 18 = $ 68

Firm should charge lower price in the market with Inelastic demand.